# Neural Bridge Processes

### v1

@@ 3.4 Bridge Construction
At test time, NBPs generate samples from the conditional distribution $p(y_{0}\mid x,\mathcal{D})$, where $\mathcal{D}=(x_{\mathbb{C}}\in\mathbb{R}^{M\times D},y_{\mathbb{C},0}\in\mathbb{R}^{M\times D})$ denotes the observed context data.
The conditional sampling proceeds as follows. First, initialize the diffusion state with the known endpoint:
$y_{T}=\gamma_{T}x+n,\quad n\sim\mathcal{N}(0,I).$ (19)
For each diffusion timestep $t=T,\ldots,1$, perform the following steps:
  • Sample the noisy context outputs using the forward diffusion bridge process:
    $y_{\mathbb{C},t}\sim\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}y_{\mathbb{C},0}+\bar{\gamma}_{t}x_{\mathbb{C}},(1-\bar{\alpha}_{t})I\right),$ (20)
  • Combine noisy target and context states at timestep $t$:
    $y_{t}=\{y_{\mathbb{T},t},y_{\mathbb{C},t}\},\quad x=\{x_{\mathbb{T}},x_{\mathbb{C}}\}.$
  • Perform the reverse diffusion step with the learned backward kernel, incorporating the bridge correction:
@@
    where
    $\mu_{\theta}(t)=\frac{1}{\sqrt{\alpha_{t}}}\left(y_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(y_{t},x,t)\right)+C_{t}(x).$ (22)
Following the Repaint Lugmayr et al. (2022) strategy, at each diffusion timestep $t$, we repeat the forward perturbation of context points and the corresponding reverse denoising step multiple times before proceeding to the next timestep. Simulating this repeated scheme from $t=T$ down to $t=1$ ensures that the context information is consistently reinforced throughout the diffusion trajectory, leading to more coherent and accurate conditional generation.
3.5 Input-output Dimensional Alignment
Real-world datasets often have mismatched input and output dimensions. To enable joint modeling in the same space, we apply a fixed projection $\mathcal{P}$ to map $x\in\mathbb{R}^{N\times D_{x}}$ to $x_{a}=\mathcal{P}(x)\in\mathbb{R}^{N\times D_{y}}$, aligning it with $y\in\mathbb{R}^{N\times D_{y}}$. For example, in image regression, $\mathcal{P}$ can add spatial or contextual information to match RGB output dimensions.
3.6 Noise Model Architecture
To ensure that our model remains consistent with the structural properties of stochastic processes and to guarantee fair experimental comparisons, we adopt the same noise model architecture as NDPs, namely the Bi-Dimensional Attention Block Dutordoir et al. (2023), as shown in Figure 1. Due to space limitations, we provide the detailed design in the Appendix.
4 Experiments
4.1 Baseline Implementation and Evaluation Metrics
To provide a comprehensive comparison, we implement NPs Garnelo et al. (2018b) , ANPs Kim et al. (2019), and ConvNPs Gordon et al. (2019) using the official NP-Family repository Dubois et al. (2020), with all hyperparameters set to the recommended default values. We also include GNP Bruinsma et al. (2021) in our synthetic experiments for comparison. For NDP Dutordoir et al. (2023), we directly utilize the official repositories. Meanwhile, to ensure a fair comparison, our NBP model adopts the same Bi-Dimensional Attention Block architecture and hyperparameters as the baseline NDP. The implementation details and additional related works are provided in the Appendix. All models are retrained on the experimental datasets to ensure consistent metric evaluation and visualization. All experiments are conducted on a single NVIDIA RTX 4090 GPU.
4.2 Regression on Synthetic Data
The log-likelihood is estimated by fitting a multivariate Gaussian to 128 samples drawn from the conditional distribution of the model. For our proposed method, we use a 4-layer transformer-style architecture with 8 attention heads and 64-dimensional hidden layers. Diffusion noise is scheduled over 500 timesteps with a cosine schedule ($\beta\in[3\mathrm{e}{-4},0.5]$). The optimizer uses a peak learning rate of $10^{-3}$ with warm-up (20 epochs) and cosine decay (200 epochs). All experiments use the same evaluation batch size and sampling procedure for consistency. For input dimensions $D=2,3$, to address the alignment issue discussed in Section 3.5, we directly define the projection operator $\mathcal{P}$ in Equation (20) as the mean of the input components. Meanwhile, the input $x$ to the denoising network $\epsilon_{\theta}$ remains unchanged.
Table 1 shows that NBP consistently outperforms prior Neural Process variants across all input dimensions. Notably, NBP maintains stable and accurate predictions in higher dimensions ($D=2,3$), where performance of other models tends to degrade sharply. Figure 2 shows representative samples generated by our model under the Squared Exponential (SE) kernel and the Matérn-5/2 kernel settings.
@@ 4.2 Regression on Synthetic Data
NBP (ours) $\mathbf{4.33}{\pm}0.03$ $\mathbf{-13.15}{\pm}0.05$ $\mathbf{-20.11}{\pm}0.04$ $0.05{\pm}0.02$ $\mathbf{-14.62}{\pm}0.03$ $\mathbf{-20.51}{\pm}0.03$
(a) Squared Exponential kernel
(b) Matérn-5/2 kernel
Figure 2: Function samples generated by NBP under two different GP kernels. In each plot, the black solid line indicates the sample mean, blue lines are function samples, red circles represent the context points, and green crosses denote the target points.
@@ 4.3.1 EEG Signal Regression Tasks
To assess NBPs on correlated multi-output prediction and missing data, we randomly mask windows in 3 of 7 channels and predict the missing values. Inputs are the concatenated temporal and channel indices $\mathbf{x}_{e}=(i_{t},i_{c})$, and outputs are the voltage measurements $\mathbf{y}_{e}$. This evaluation is carried out under three distinct experimental settings:
  • Interpolation: Predicting missing values within the existing temporal span.
  • Reconstruction: Predicting values from partially obscured temporal segments.
  • Forecasting: Predicting future values beyond the observed temporal data.
Performance metrics employed include Mean Squared Error (MSE) and Negative Log-Likelihood (NLL). As demonstrated in Table 2, the NBP consistently surpasses baseline methods across all three scenarios, underscoring NBPs’ proficiency in capturing intricate temporal structures and cross-channel dependencies inherent in EEG data.
Table 2: Predictive NLL ($\downarrow$) and MSE ($\downarrow$) on EEG
Method Inter. Recon. Forec.
NLL MSE($\times 10^{-2}$) NLL MSE($\times 10^{-2}$) NLL MSE($\times 10^{-2}$)
NP 1.66 0.52 1.78 0.44 1.61 0.39
ANP 0.47 0.25 0.70 0.48 0.90 0.60
ConvNP 0.44 0.40 -2.43 0.40 -2.34 0.55
NDP -2.46 0.18 -2.59 0.23 -2.69 0.38
NBP (Ours) -3.35 0.16 -3.22 0.16 -3.51 0.29
(a) Ground Truth
(b) Corrupted Context
(c) Model Prediction
Figure 3: Qualitative Results on CelebA $\times64$ Image Regression Task: (a) Ground Truth, (b) Corrupted Context, (c) Model Prediction. The task requires inferring the true image content based on randomly corrupted context information and the spatial coordinates of target pixels.
@@ 4.3.2 Image Regression Task
Figure 4: Comparison of NBP and NDP: NBPs consistently achieve lower loss values across the majority of training iterations
@@
In this experiment, we apply Neural Bridge Processes (NBPs) to the image regression task, where the objective is to predict pixel values based solely on their spatial coordinates normalized within the range [2,2][-2,2]. We conduct experiments using the CelebA dataset at resolutions of $32\times 32$ and $64\times 64$. The experimental setup, including hyperparameters, denoising network architecture, learning rate, random seed, and other baseline configurations, strictly follows the standard configuration used by NDPs for a fair comparison.
We evaluate the performance of NBPs under different levels of context sparsity and directly compare their predictive accuracy with that of NDPs. The Mean Squared Error (MSE) results are summarized in Figure 5, with detailed numerical values provided in Tables 4 and 5 in the Appendix. As shown in Figure 5, the proposed NBP (orange) consistently achieves lower MSE across all context ratios compared to the NDP baseline (blue). Moreover, NBPs exhibit smaller standard deviations, indicating more stable and reliable predictions. All MSE values are computed by averaging over nine conditional samples for each input, with pixel values normalized to the range $[0,1]$. Figure 3 provides an illustrative depiction of the image regression task setup.
Under various observation conditions, NBPs consistently outperform NDPs by a substantial margin. For example, at a context ratio of $0.02$, NBPs achieve an MSE of $0.76$ compared to $0.88$ by NDPs on CelebA $32\times 32$. This demonstrates NBPs’ enhanced capability to capture spatial dependencies and deliver accurate predictions even from highly sparse context observations. This advantage extends to higher resolutions: at the same context ratio ($0.02$) on CelebA $64\times 64$, NBPs achieve an MSE of $0.80$ compared to NDPs’ $1.05$, underscoring the scalability and robustness of our method.
Furthermore, Figure 4 visualizes the denoising score matching objective loss during training, illustrating that NBPs consistently achieve lower loss values across the majority of training iterations. This result supports the conclusion that the bridge-based training paradigm significantly enhances the effectiveness of denoising diffusion probabilistic model (DDPM) path supervision.
@@ 4.4 Computational Efficiency
We observe in our experiments that, under consistent training settings—including the denoising network architecture, learning rate, random seed, and other base hyperparameters—the proposed NBP model and the baseline NDP consume approximately the same amount of time per epoch and in total. This indicates that NBP does not introduce additional computational overhead.
This efficiency stems from the fact that NBP does not incorporate any extra architectural complexity. Instead, it enhances the training signal through a coupling mechanism between the inputs and outputs in the neural diffusion process (as described in Section 3). This coupling is implemented entirely at the software level, without increasing the model’s structural depth or parameter count.
@@ 5 Conclusion
In this work, we introduced Neural Bridge Processes (NBPs), a diffusion-based framework for stochastic function modeling that explicitly incorporates input supervision throughout the diffusion trajectory. By reformulating the forward kernel with a principled bridge coefficient, NBPs address the weak input coupling and endpoint mismatch of traditional Neural Diffusion Processes (NDPs), ensuring stronger conditional guidance and better theoretical consistency. Unlike computationally intensive SDE-based bridges, NBPs implement bridge corrections efficiently within the DDPM framework using SNR- and path-aware modeling. Experiments on synthetic data, real-world EEG time series, and image regression tasks demonstrate that NBPs significantly enhance predictive accuracy and uncertainty calibration compared to state-of-the-art NDP baselines. These results highlight the potential of NBPs for structured generative modeling, paving the way for future extensions to high-dimensional, multi-modal, and control-oriented applications.
Appendix A Related Works

### v2

@@ 4.3 Bridge Construction
At test time, NBPs sample from $p(y_{0}\mid x,\mathcal{D})$, where $\mathcal{D}=(x_{\mathcal{C}},y_{\mathcal{C},0})$ denotes the observed context set. We initialize the terminal state from the input-aligned bridge endpoint,
$y_{T}=\bar{\gamma}_{T}x+n,\quad n\sim\mathcal{N}(0,(1-\bar{\alpha}_{T})I).$ (15)
For each reverse timestep, the observed context outputs are re-noised using the closed-form bridge marginal
$y_{\mathcal{C},t}\sim\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}y_{\mathcal{C},0}+\bar{\gamma}_{t}a_{\psi}(x_{\mathcal{C}}),(1-\bar{\alpha}_{t})I\right),$ (16)
then combined with the current target state as $y_{t}=\{y_{\mathcal{T},t},y_{\mathcal{C},t}\}$ and $x=\{x_{\mathcal{T}},x_{\mathcal{C}}\}$. The reverse transition is then sampled from
@@
where $\mu_{\theta}$ includes the bridge correction in Eq. (13). Following RePaint Lugmayr et al. (2022), we may repeat the context re-noising and reverse denoising step at each timestep before moving to t1t-1. This repeatedly reinjects the observed context and improves conditional coherence along the generated trajectory.
4.4 Theoretical Analysis: Why Process-Level Anchoring Helps
The preceding propositions establish that NBP is a valid DDPM-style bridge construction. We next make explicit what is gained by placing the input in the forward process rather than only in the denoising network. The goal is not to claim a universal generalization guarantee, but to identify structural effects that follow directly from the bridge kernel.
Theorem 4.4 (Pathwise input identifiability).
Fix $y_{0}$ and consider two inputs $x,x^{\prime}$. Let $q_{t}^{\mathrm{NBP}}(\cdot|y_{0},x)$ denote the NBP forward marginal in Eq. (8). Then
$W_{2}\!\left(q_{t}^{\mathrm{NBP}}(\cdot|y_{0},x),q_{t}^{\mathrm{NBP}}(\cdot|y_{0},x^{\prime})\right)=|\bar{\gamma}_{t}|\,\|x-x^{\prime}\|_{2}.$ (18)
By contrast, the NDP forward marginal is independent of $x$, and therefore the corresponding Wasserstein distance is zero for all $x,x^{\prime}$.
Theorem 4.4 formalizes the sense in which NBP creates pathwise input coupling: different inputs induce different noisy states along the forward trajectory. A larger conditioning network can make the reverse model more expressive, but it does not change the fact that the NDP training states are generated from an input-independent forward kernel.
Theorem 4.5 (Forward-state information about the input).
Assume $x\sim\mathcal{N}(0,\Sigma_{x})$ and condition on $y_{0}$. Under the NBP forward marginal,
$I(x;y_{t}\mid y_{0})=\frac{1}{2}\log\det\!\left(I+\frac{\bar{\gamma}_{t}^{2}}{1-\bar{\alpha}_{t}}\Sigma_{x}\right).$ (19)
For the NDP forward process, $I(x;y_{t}\mid y_{0})=0$ because $y_{t}$ is conditionally independent of $x$ given $y_{0}$.
This gives an information-theoretic interpretation of the bridge term. NBP makes the noisy training state itself informative about the input; NDP leaves all input information to the denoiser-side conditioning pathway. The amount of information depends on the schedule through the signal-to-noise ratio $\bar{\gamma}_{t}^{2}/(1-\bar{\alpha}_{t})$, which also motivates schedule selection.
Proposition 4.6 (Direct gradient pathway).
For a single-timestep denoising loss
$\ell_{t}(\theta,x)$ $=\|\epsilon_{\theta}(y_{t},x,t)-\epsilon\|_{2}^{2},$ (20)
$y_{t}$ $=\sqrt{\bar{\alpha}_{t}}y_{0}+\bar{\gamma}_{t}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon,$
the total derivative with respect to $x$ decomposes as
$\frac{d\ell_{t}}{dx}=\underbrace{\frac{\partial\ell_{t}}{\partial x}}_{\text{denoiser conditioning path}}+\underbrace{\bar{\gamma}_{t}\frac{\partial\ell_{t}}{\partial y_{t}}}_{\text{forward bridge path}}.$ (21)
For NDP, $\bar{\gamma}_{t}=0$ in the forward process, so only the denoiser conditioning path remains.
Proposition 4.6 explains why the bridge construction can improve training even when the denoising architecture is unchanged. NBP supplies an additional route by which input supervision affects the denoising loss through the sampled state $y_{t}$ itself. This complements the empirical Jacobian sensitivity in Figure 1, where NBP shows stronger input sensitivity especially in early diffusion steps.
@@ 4.4 Theoretical Analysis: Why Process-Level Anchoring Helps
SNP $4.27{\pm}0.02$ $13.19{\pm}0.03$ $20.24{\pm}0.04$ $0.01{\pm}0.02$ $14.67{\pm}0.03$ $20.59{\pm}0.05$
GEOMNDP $4.22{\pm}0.05$ $13.36{\pm}0.05$ $20.45{\pm}0.05$ $0.13{\pm}0.03$ $14.73{\pm}0.03$ $20.63{\pm}0.06$
NBP (ours) $\mathbf{4.72}{\pm}0.03$ $\mathbf{-12.10}{\pm}0.05$ $\mathbf{-18.10}{\pm}0.04$ $0.047{\pm}0.019$ $\mathbf{-13.30}{\pm}0.03$ $\mathbf{-18.46}{\pm}0.03$
Table 2: Predictive NLL ($\downarrow$) and MSE ($\downarrow$) on EEG
Method Interpolation Reconstruction Forecasting
NLL MSE($\times 10^{-2}$) NLL MSE($\times 10^{-2}$) NLL MSE($\times 10^{-2}$)
NP $1.66\pm 0.08$ $0.52\pm 0.03$ $1.78\pm 0.09$ $0.44\pm 0.02$ $1.61\pm 0.07$ $0.39\pm 0.02$
ANP $0.47\pm 0.05$ $0.25\pm 0.02$ $0.70\pm 0.06$ $0.48\pm 0.03$ $0.90\pm 0.06$ $0.60\pm 0.04$
ConvNP $0.44\pm 0.04$ $0.40\pm 0.03$ $2.43\pm 0.12$ $0.40\pm 0.02$ $2.34\pm 0.11$ $0.55\pm 0.03$
NDP $2.46\pm 0.11$ $0.18\pm 0.01$ $2.59\pm 0.13$ $0.23\pm 0.02$ $2.69\pm 0.12$ $0.38\pm 0.02$
SNP $3.19\pm 0.10$ $0.16\pm 0.01$ $\mathbf{-3.30\pm 0.11}$ $0.18\pm 0.01$ $3.02\pm 0.09$ $0.31\pm 0.02$
GEOMNDP $2.48\pm 0.12$ $0.18\pm 0.01$ $2.65\pm 0.13$ $0.20\pm 0.01$ $2.84\pm 0.11$ $0.34\pm 0.02$
NBP (Ours) $\mathbf{-3.55\pm 0.10}$ $\mathbf{0.15\pm 0.01}$ $\mathbf{-3.45\pm 0.11}$ $\mathbf{0.15\pm 0.01}$ $\mathbf{-3.74\pm 0.11}$ $\mathbf{0.27\pm 0.02}$
5 Experiments
5.1 Baseline Implementation and Evaluation Metrics
For a comprehensive comparison, we implement Neural Processes (NPs) Garnelo et al. (2018b), Attentive Neural Processes (ANPs) Kim et al. (2019), and Convolutional Neural Processes (ConvNPs) Gordon et al. (2019) using the official NP-Family repository Dubois et al. (2020), with all hyperparameters set to the recommended defaults. We further include Gaussian Neural Processes (GNPs) Bruinsma et al. (2021) in our synthetic experiments. For Neural Diffusion Processes (NDPs) Dutordoir et al. (2023), Geometric Neural Diffusion Processes (GEOMNDPs) Mathieu et al. (2023), and Score-Based Neural Processes (SNPs) Dou et al. (2025), we directly adopt their official implementations. To ensure a controlled comparison with NDP, our Neural Bridge Process (NBP) uses the same Bi-Dimensional Attention Block denoising backbone, optimizer, diffusion schedule, number of denoising steps, and training budget as the NDP baseline. The bridge coefficients and reverse correction introduce no trainable parameters. When $D_{x}\neq D_{y}$, NBP additionally uses the learned anchor map $P_{\psi}$, which is part of the proposed bridge construction rather than a replacement for the denoiser; we therefore report fixed-anchor and learned-anchor ablations in Appendix H.2 to separate this effect from the bridge itself. Detailed implementation settings are provided in Appendix G. Additional ablation studies, including bridge coefficient schedules, component ablations on CelebA NLL, a bridge-transfer experiment on FlowNP, and an analysis quantifying path supervision, are presented in Appendices H, H.2, and H.3. All models are retrained on the respective experimental datasets and evaluated with the same set of random seeds; reported values are averages over multiple seeds unless otherwise specified. Experiments are conducted on NVIDIA RTX A6000 GPUs with 48GB memory; each reported run uses a single GPU unless otherwise stated.
5.2 Regression on Synthetic Data
The log-likelihood is estimated by fitting a multivariate Gaussian to 128 samples drawn from the conditional distribution of the model. For our proposed method, we use a 4-layer transformer-style architecture with 8 attention heads and 64-dimensional hidden layers. Diffusion noise is scheduled over 500 timesteps with a cosine schedule ($\beta\in[3\mathrm{e}{-4},0.5]$). The optimizer uses a peak learning rate of $10^{-3}$ with warm-up (20 epochs) and cosine decay (200 epochs). All experiments use the same evaluation batch size and sampling procedure for consistency. When input and output dimensions differ, the bridge uses the learned output-space anchor $a_{\psi}(x)=P_{\psi}(x)$ described in Appendix D; the raw input $x$ is still provided to the denoising network $\epsilon_{\theta}$ as conditioning information.
Table 1 shows that NBP consistently outperforms prior Neural Process variants across all input dimensions. Notably, NBP maintains stable and accurate predictions in higher dimensions ($D=2,3$), where performance of other models tends to degrade sharply.
@@ 5.3.1 EEG Signal Regression Tasks
To assess NBPs on correlated multi-output prediction and missing data, we randomly mask windows in 3 of 7 channels and predict the missing values. Inputs are represented by the concatenated temporal and channel indices $\mathbf{x}_{e}=(i_{t},i_{c})$, and outputs are the corresponding voltage measurements $\mathbf{y}_{e}$. We evaluate three conditional prediction regimes in a unified setting. In interpolation, the model recovers locally missing values within the observed temporal span. In reconstruction, the model infers partially obscured temporal segments of a target channel using the remaining observed channels as context. In forecasting, the model extrapolates future signal trajectories from current observations. These regimes test complementary aspects of conditional function modeling: local completion, cross-channel reconstruction, and temporal extrapolation. We report Mean Squared Error (MSE) and Negative Log-Likelihood (NLL). As shown in Table 2, NBP consistently outperforms the baselines across the three settings, indicating that the bridge-anchored diffusion process can capture both temporal structure and cross-channel dependencies in EEG signals. Appendix Table 3 further reports calibration metrics on the interpolation setting, where NBP achieves the lowest ECE and sharpest predictive uncertainty while maintaining coverage close to the nominal 90% level.
CylinderFlow.
To further evaluate the empirical scope of NBP beyond synthetic functions and EEG signals, we also report results on the CylinderFlow benchmark, which requires structured spatial prediction under partial observations. Table 5 shows that NBP achieves the lowest MSE among all compared methods, improving over both classical NP variants and the diffusion-based NDP baseline. This suggests that the benefit of bridge-anchored diffusion generalizes to real-world spatial prediction tasks.
Table 4: MSE ($\downarrow$) on the CylinderFlow benchmark.
Method MSE
NP 0.0108
ANP 0.0157
ConvNP 0.0341
NDP 0.0053
NBP (Ours) 0.0031
Table 5: CelebA image regression NLL ($\downarrow$) under half-image and random-context observations.
Method Half Random
NP -0.84 -1.07
ANP -1.39 -1.43
ConvNP 0.11 -1.67
NDP -3.92 -3.98
SNP -4.76 -4.77
NBP (Ours) -5.02 -5.24
@@ 5.3.2 Image Regression Task
We evaluate image regression on CelebA at resolutions of $32\times 32$ and $64\times 64$, where models predict pixel values from spatial coordinates normalized to [2,2][-2,2]. Hyperparameters, denoising architecture, learning rate, seed set, and baseline configurations follow the NDP setup for a controlled comparison. Our baselines focus on NP-family and diffusion/score-based Neural Process models, which share the same conditional function-learning setting.
Figure 2 summarizes MSE across context ratios, with numerical values in Appendix Tables 7 and 8 and qualitative examples in Appendix Figure 6. NBP consistently improves over the NDP baseline and has smaller standard deviations across context ratios. All MSE values average nine conditional samples per input, with pixel values normalized to $[0,1]$. Table 5 further shows NLL improvements over broader NP-family baselines, while Appendix H.2 decomposes the gain into endpoint initialization, bridge construction, anchor learning, and denoiser-capacity effects.
For example, at context ratio $0.2$, NBP reduces MSE from $0.88$ to $0.76$ on CelebA $32\times 32$ and from $1.05$ to $0.86$ on CelebA $64\times 64$, supporting the usefulness of bridge-anchored trajectories for sparse-coordinate image regression.
@@ 5.4 Computational Efficiency
Under identical training settings, NBP and the NDP baseline exhibit comparable training time per epoch and overall runtime. NBP uses the same denoising architecture and number of sampling steps as NDP; the additional operations are limited to schedule-dependent bridge coefficients and a simple correction term. Thus, the practical overhead is negligible in our implementation, while the improvement comes from the process-level input–output coupling rather than additional network capacity.
@@ 6 Conclusion
In this work, we introduced Neural Bridge Processes (NBPs), a diffusion-based framework for stochastic function modeling that explicitly incorporates input supervision throughout the diffusion trajectory. By reformulating the forward kernel with a bridge coefficient and, when needed, a learned output-space anchor, NBPs address the weak input coupling and endpoint mismatch of traditional Neural Diffusion Processes (NDPs) while preserving the same denoising backbone. Experiments on synthetic data, EEG time series, CylinderFlow, and image regression show improved predictive accuracy and uncertainty calibration. The component ablations and FlowNP transfer study further suggest that the benefit comes from input-anchored generative paths rather than only additional capacity or a DDPM-specific implementation. These results highlight a broader possibility: bridge-anchored paths may be a useful design principle for generative Neural Processes, including diffusion-, score-, and flow-based variants.
Appendix A Additional Related Discussion